OpenStudy (anonymous):

|2z-7|=13 Find the solution

OpenStudy (anonymous):

A. 10 B. -3, 10 C. -6, 10 D. No solution

OpenStudy (shamil98):

\[|2z-7| = 13 \] \[2z-7=13\] \[-(2z-7)=13 ,or 2z-7 = -13\]

OpenStudy (shamil98):

solve for z.

OpenStudy (anonymous):

-3? @shamil98

OpenStudy (shamil98):

2z-7 = 13 2z = 20 z= 10 2z- 7 = -13 2z = -6 z= -3 so yes -3 and 10,

OpenStudy (anonymous):

Thank you!